# suppSupplement

### v1

@@ Abstract
We apply this extended optimization scheme to discover the simplest setup to implement an optomechanical circulator in the experimental scenario considered in [40]. Three microwave port modes are coupled via the optomechanical interactions mediated by a varying number of mechanical modes. No direct coupling is available (see Fig. 4(a)). Since each optomechanical coupling requires a different laser drive, the key measure of experimental complexity in this case is the overall number of couplings. The hand-crafted architecture proposed in [40] requires two far-detuned mechanical modes and six couplings. Our optimization scheme finds a simpler setup, which requires the same number of far-detuned modes but only five couplings (see Fig. 4(b)). If we apply the softer constraint where passive couplings are allowed between the optical modes, our scheme identifies even simpler graphs, only involving one mechanical mode (see Fig. 4(c)). To reach the ideal behaviour, e.g. , for the left graph in Fig. 4(c), the cooperativities $C_{ab}$ and $C_{bb}$ have to tend towards infinity, while obeying that the ratio $C_{ab}^{2}/C_{bb}$ equals $1$ (see Fig. 4(d)).
@@
In conclusion, we have presented an automated-discovery approach that finds irreducible scattering setups, leading to results that provide new insights and are interpretable and generalizable. The approach is amenable to numerous future extensions like search for optimized implementations in specific hardware platforms neural-network-guided heuristics for larger devices where complete enumeration is infeasible, or target specific quantum noise characteristics. More generally, our results now open the door towards automated discovery in additional related domains of high importance, like transport through periodic structures composed of small modular blocks or scattering-based sensing experiments. Nonlinear setups or even devices with functionalities based on time-dependent control require significant new developments, but still could be treated with the same graph-based efficient search algorithm. Overall, we feel that the results shown here validate the fruitfulness of the main ideas guiding the field of artificial scientific discovery.
@@ Acknowledgements.
We acknowledge fruitful discussions with Mario Krenn, Clara Wanjura, and Flore Kunst.
Appendix A Details on the high-level description of a multi-mode circuit
Here, we give more details on our high-level description of multi-mode circuits. For completeness, we consider the most general scenario which encompasses purely-routing devices as well as phase-preserving, and phase-sensitive amplifiers.
A.1 Bogoliubov de Gennes Hamiltonian and the rotating wave approximation
@@
The Hamiltonian Eq. 3 is defined in a frame in which the phase space of each mode $j$ rotates at the corresponding carrier frequency $\omega_{L,j}$. Thus, $-g_{jj}$ is to be interpreted as the detuning $\Delta_{j}$ of the carrier frequency $\omega_{L,j}$ from the mode resonance $\omega_{j}$, $\Delta_{j}=\omega_{L,j}-\omega_{j}$. For multi-mode circuits, such time-independent Hamiltonian is typically referred to as the rotating-wave Hamiltonian as it is obtained after a Rotating-Wave Approximation (RWA) that drops the fast-oscillating terms, which result from non-resonant interactions. The corrections to the RWA approximation are small if all resonant frequencies $\omega_{j}$ are much larger than the coupling rates $g_{ij}$ and $\nu_{ij}$ as well as the decay rates $\kappa_{j}$. We note that if all couplings are active the carrier frequencies $\omega_{L,j}$ for the different modes can be chosen independently of each other, as long as the frequencies of the lasers mediating the interactions are adjusted accordingly. Alternatively, one can use this freedom to choose independently the values of the detunings $\Delta_{j}$ in experimental scenarios in which the mode resonances $\omega_{j}$ are not tunable. We note that all modes connected by passive coupling have the same carrier frequency which in turn leads to constraints to $\Delta_{j}$ for fixed $\omega_{j}$. Besides these constraints, the mode resonances $\omega_{j}$ are irrelevant in the high-level description based on the RWA. This is an appealing feature that makes any insight gained using our method highly transferable within a variety of platforms. A more refined description including non-resonant interactions is typically used to quantify how a specific implementation deviates from the ideal target behavior. This deviation is device-specific and, thus, goes beyond the scope of our work.
A.2 Langevin Equations
@@
$\dot{\hat{a}}_{j}(t)$ $=-\mathrm{i}\sum_{k}g_{jk}\hat{a}_{k}(t)-\mathrm{i}\sum_{k}\nu_{jk}\hat{a}_{k}^{\dagger}(t)-\frac{\kappa_{j}+\Gamma_{j}}{2}\hat{a}_{j}(t)-\sqrt{\kappa}_{j}\hat{a}_{j}^{\mathrm{in}}(t)-\sqrt{\Gamma}_{j}\hat{a}_{j}^{\mathrm{noise}}(t)$ (5)
Here, $\hat{a}_{j}^{\mathrm{in}}$ is the input field for the main decay channel (for the port modes the out-coupling channel), with decay rate $\kappa_{j}$. Likewise, $\hat{a}_{j}^{\mathrm{noise}}$ is the noise entering a port mode $j$ from its intrinsic loss channel, which has decay rate $\Gamma_{j}$. The Langevin equations can be rewritten in a compact form as
$\dot{\hat{\boldsymbol{\xi}}}(t)$ $=\left(-i\sigma_{z}H_{\rm BdG}-\frac{\kappa+\Gamma}{2}\right)\hat{\boldsymbol{\xi}}(t)-\sqrt{\kappa}\hat{\boldsymbol{\xi}}^{\mathrm{in}}(t)-\sqrt{\Gamma}\hat{\boldsymbol{\xi}}^{\mathrm{noise}}(t)$ (6)

### v2

@@ VI Discovering asymptotic solutions
We apply this extended optimization scheme to discover the simplest setup implementing an optomechanical circulator in the experimental scenario considered in [41]. Three microwave port modes are coupled via the optomechanical interactions mediated by a varying number of mechanical modes. No direct coupling is available (see Fig. 6(a)). Since each optomechanical coupling requires a different laser drive, the key measure of experimental complexity in this case is the overall number of couplings. The hand-crafted architecture proposed in [41] requires two far-detuned mechanical modes and six couplings (see Fig. 6(b)). Our optimization scheme finds a simpler setup, which requires the same number of far-detuned modes but only five couplings (see Fig. 6(c)). If we apply the softer constraint where passive couplings are allowed between the optical modes, our scheme identifies even simpler graphs, only involving one far-detuned mode (see Fig. 6(c)).
As we explained above, Autoscatter translates the asymptotic graphs with far-detuned modes into an effective graph featuring only port modes. Our effective graphs improve the interpretability of the solutions. For example, all valid asymptotic graphs in our optomechanical circulator are translated into the same effective three-mode circulator graph shown in Fig. 6(d), see Supplemental Material E for details. As we discussed in Section III, this graph encodes the most general three-modes solution for an ideal circulator/isolator, see Eq. 4. We can leverage this solution to discover the manifold of asymptotic parameters corresponding to ideal solutions for our discovered asymptotic graphs. As an example, consider the lower left graph in Fig. 6(c). Here, all passive couplings between the optical modes would be allowed. However, they can not realize the required synthetic field flux (see Section VIII). This role is fulfilled by the mechanical mode, which mediates an effective complex-valued beamsplitter coupling between $\hat{a}_{1}$ and $\hat{a}_{2}$. We empirically find that this architecture has only solutions where the effective detuning is zero. Therefore, the effective cooperativity $C_{a_{1}a_{2}}^{\mathrm{eff}}$ should fulfill $C_{a_{1}{a_{2}}}^{\mathrm{eff}}=1$, see Eq. (4). Taking into account that $C_{a_{1}a_{2}}^{\mathrm{eff}}=C_{a_{1}b}C_{a_{2}b}/C_{bb}$, this constraint identifies a manifold of ideal solutions for the asymptotic parameters $C_{a_{1}b}/\sqrt{C_{bb}}$ and $C_{a_{2}b}/\sqrt{C_{bb}}$, as visualized in Fig. 6(e).
In realistic experimental scenarios, the coupling strength and loss rates have upper and lower bounds limiting how close one can approach the asymptotic limits, see Fig. 6(f). For example, for the lower left graph in Fig. 6(c), the loss function $\cal{L}$ scales like ${\cal L}\sim r/C_{bb}$ in the asymptotic limit. Here, $r$ is a constant that can be interpreted as a convergence speed. Even though in the asymptotic limit, all solutions are ideal, they differ in the convergence speed. In our example, when we send the dimensionless detuning $C_{bb}$ to infinity, the convergence speed $r$ depends on the ratio of the cooperativities, see inset in Fig. 6(e). For practical applications, similar considerations can be used to identify a preferred path to approach the asymptotic limit.
VII Analysis of the search space reduction
Figure 7: Search space reduction. We compare the number of graphs tested by AutoScatter to the total number of graphs within the search space. A brute-force search would have to test every graph (dashed black line). Each data point corresponds to a different target functionality/device: (i) Gyrator, (ii) Isolator, (iii) Amplifier with intrinsic port losses, (iv) Three-port gyrator, (v) Wilkinson power divider, (vi) Amplifier without intrinsic port losses, (vii) Circulator with only passive couplings between optical modes, (viii) Directional coupler, (ix) Resistive power divider, and (x) Circulator without couplings between optical modes. All examples marked with an asterisk are discussed in the Sections III, IV, V and VI. See Supplemental Material G for a summary of all further examples.
The main achievement of our work is the automated discovery of interpretable and transferable coupled mode setups. In addition, we have introduced an efficient discrete optimization algorithm that allows us to reduce the search space compared to a brute-force search. An interesting open question is how large the reduction in the search space will be. Considering that our algorithm can be directly applied to solve a large class of problems whose potential solutions can be encoded in graphs with continuous weights, this question is of interest beyond our setting of coupled mode setups.
In Fig. 7, we show the number of graphs tested by AutoScatter in comparison to the total number of graphs within the search space. Each data point corresponds to a different target device, as identified by the target scattering matrix.
As we can see, the performance gain in comparison to a brute-force approach was highly significant in some of the most complex problems. If we quantify the performance gain as the ratio of the number of graphs in the search space over the number of tested graphs, it varied in a broad range from $\sim 5$ (see, e.g., (vii)) to $\sim 400$ in our experiments (see (ix)). As one might intuitively expect, the largest gains are obtained for complex devices featuring four or five modes. This is because the brute-force search space grows exponentially with the squared number of modes in the device. It is not possible to predict in general how the performance gain evolves with mode number, since that depends on the target device. In Supplemental Material F, we provide more details on our analysis and discuss how the search space structures of the individual target devices affect the respective performance gain. Furthermore, we provide in Supplemental Material G a summary of all target devices included in Fig. 7. With the help of the substantial performance gains displayed here, and for the mode numbers relevant for the variety of practically important target devices, we found our approach to be able to provide the complete enumeration of solutions that is one of its key advantages. It should be noted that eventually, for even further increased mode numbers, one would have to resort to a combination of our search-space reduction with heuristic approaches.
VIII Implementation in specific experimental platforms
[…]
Typically, the first concrete choice in view of designing a scattering device is the type of port modes. This choice will depend on the functionality of the target device, e.g., for a microwave to optics converter, one needs a microwave and an optical mode with resonance frequency close to the respective target carrier frequencies.
Pair of modes of the same type and with similar resonance frequency can naturally be coupled via a passive beamsplitter interaction [42, 8], e.g., the evanescent coupling between two defect modes in a photonic crystal [8]. By itself, even a large difference between the resonance frequencies of two modes does not pose a fundamental constraint to their available linearized couplings because signals can be up and down converted by pumping a weak nonlinearity with a coherent drive. The specific coupling scheme depends on the experimental platform. In a purely optical setting, one can exploit the optical nonlinearities in the material, i.e., the Kerr or the $\chi^{(2)}$ nonlinearity. Mechanical modes and optical modes can be coupled via radiation pressure. In superconducting circuits, the coupling of microwave modes of different frequencies is based on the Josephson nonlinearity, driving the circuit with an AC current or a time-dependent magnetic flux piercing a Superconducting Quantum Interference Device (SQUID) [49]. While active couplings are widely available, some constraints to the quadratic Hamiltonian Eq. 1 and, thus, our graphs connectivity might be present in specific experimental settings. For example, coupling a circuit microwave mode to an optical cavity mode typically requires a mechanical auxiliary mode [5].
The next and most important step in designing a new scattering experiment is to find a suitable RWA Hamiltonian that enforces the desired scattering behavior, taking into account possible hardware-specific constraints on the couplings. This step typically requires human ingenuity and is the central focus of many experimental and theoretical works [3, 42, 6, 8, 31, 32, 36, 37, 39, 41]. Our method fully automates this task finding all possible solutions and representing them using our graph language (see Fig. 1(a)).
At the level of the RWA, the solutions are highly transferable between hardware platforms as one needs only to commit to a certain number of modes, coupling types, and the underlying cooperativites and synthetic field fluxes, as we have shown in Appendix F. Since they do not depend on the resonance frequency of the modes, they leave open the choice of whether to implement certain couplings as active or passive couplings. For example, the isolator graph shown in Fig. 4(a) can be implemented fully actively, requiring three parametric pumps, as in [2]. Alternatively, it can be implemented using two parametric pumps and a passive coupling as in [42].
If one is willing to employ active couplings for every link, any desired graph can certainly be implemented. This approach, combining several pump tones at different frequencies, is common in different platforms [37, 41, 3, 5, 38, 33, 1, 2, 4, 7]. However, implementing active couplings that rely on parametric driving may not be the most economical design in many scenarios. For example, in the optical regime, it is more challenging to combine many pump tones at different frequencies in this higher frequency range. For this reason, it is valuable to identify those setups that require the minimum amount of parametric pumps.
Figure 8: Minimization of the number of active couplings. The graphs are annotated with the carrier frequency classes that minimize the number of active couplings for the directional coupler graphs from Fig. 4(b). Edges that connect two modes belonging to the same frequency class are coupled passively. The remaining couplings are active. A sorting is correct when it fulfills the constraint that complex-valued beamsplitter couplings (green) and squeezing (blue) interactions have to connect modes of different frequency classes. We find all possible sortings by using a brute-force search, see Appendix I for more details. The shown sortings minimize the number of active couplings. Graph (I) has six more such solutions (not shown).
In many experimental platforms, one needs at least one active coupling to engineer a synthetic gauge flux [37, 41, 3, 38, 33, 1, 2, 4] , $\Phi_{i,j,l,\ldots,k}=\arg(g_{ij}g_{jl}\ldots g_{ki})\neq 0$. A notable exception are setups featuring passive coupling between chiral modes [50, 51], see Supplemental Material B. Here, we exclude such exceptions and assume that all loops formed exclusively by passive couplings have zero flux. In this scenario, we can always choose a gauge in which all passive couplings are real-valued. However, only beamsplitter interactions coupling modes with the same carrier frequency can be realised passively, and this could be forbidden by other constraints in the graph. In particular, all complex-valued beamsplitter couplings and squeezing couplings are active and, thus, connect modes of different frequencies.
To determine the minimum number of active couplings, one has to solve a sorting problem, in which one sorts modes into frequency classes. If two modes are in the same class, they have the same carrier frequency, otherwise, they do not. The goal is to minimize the number of edges connecting modes of different classes under the aforementioned constraints. We solve this problem with a brute-force search, see Appendix I for more details. This search yields the minimum number of pumps as well as all possible combinations of passively coupled modes that allow to reach the minimum. As an example, we analyze the number of active pumps for the fully directional coupler of Fig. 4(b). Our analysis shows that Graph (I) requires at least four parametric pumps, while graph (II) and (III) only require three pumps. Fig. 8 displays two different optimal solutions for each graph. Graphs (II) and (III) have only these two solutions, while graph (I) has eight in total.
Figure 9: Frequency-dependent scattering response. We show the scattering response of a directional coupler (Graph (1) in Fig. 4(b)) for different values of the loss rates. (a) Graph representation labeled with the out-couplings (orange arrows) and intrinsic losses (red arrows). (b) Transmission from an input to the output port (top) and reverse transmission (bottom) as a function of the frequency (counted off from the carrier frequency) for different outcoupling ratios $\kappa_{\mathrm{in}}/\kappa_{\mathrm{out}}$ (see colorbar), and $\kappa_{\mathrm{aux}}/\kappa_{\mathrm{out}}=2$ (solid lines) and $10$ (dashed lines). The ideal transmission is |$\absolutevalue{S_{\mathrm{out,in}}}=1/\sqrt{2}$ (dashed black line) while the ideal reverse transmission $S_{\mathrm{in,out}}$ is zero. Our approach optimised the system, such that the coupler always displays the ideal behaviour at the carrier frequency (for $\omega=0$). (c) Bandwidth as function of $\kappa_{\mathrm{in}}$ for $\kappa_{\mathrm{aux}}/\kappa_{\mathrm{out}}$ as in (b). We define the bandwidth as the full width at half maximum (FWHM) of the transmission |$\absolutevalue{S_{\mathrm{out,in}}}$ |.
Even after choosing a graph and deciding which couplings are active and passive, our solutions still leave the freedom to choose the modes’ decay rates. We recall that the weights $H_{ij}$ of our graphs are proportional to square roots of cooperativities, $H_{ij}=g_{ij}/\sqrt{\kappa_{i}\kappa_{j}}$ or $H_{ij}=\nu_{ij}/\sqrt{\kappa_{i}\kappa_{j}}$ cf. Eq. 2. Independent of the concrete values for the modes’ decay rates $\kappa_{i}$, the scattering amplitudes at the carrier frequencies do not change as long as the coupling rates $g_{ij}$ and $\nu_{ij}$ are rescaled accordingly. At the same time, the outcouplings and other loss rates can typically be tuned to a certain degree in experiments. This freedom can be used to change the frequency-dependence of the scattering process, as illustrated in Fig. 9 for one of our directional couplers. By increasing the outcoupling rate of the input ports, one can increase the bandwidth of the transmission from the input to the output. With large losses on the auxiliary modes, one can suppress the reverse transmission over a broad frequency range.
Beyond the design choices discussed above, any hardware platform possesses its own hardware-specific challenges. The device should be engineered to host the desired high-quality-factor modes with the target passive quadratic couplings and strong enough high-order nonlinear interactions. This task requires domain-specific simulations of, e.g., the Maxwell equations or the elasticity equations. We note in passing that this task typically requires human ingenuity and, for this reason, has been the focus of intense efforts aimed at automatizing it, see e.g. [52, 53, 54]. However, its platform-specific character makes any finding less transferable and generalizable. For this reason, it is outside of the scope of our work.
As a final step, the dynamics of the multimode system can be modelled using a set of coupled Langevin equations featuring also the nonlinear interactions, with bare couplings calculated from the domain-specific simulations. One can straightforwardly use this more general description to find the driving protocol to implement the target active couplings as detailed below. The pump frequencies $\omega_{ij}^{P}$ are chosen to up and down-convert the signals to the target carrier frequencies $\omega_{L,j}$, e.g. for a nonlinearity based on three-wave mixing $\omega_{ij}^{P}=\absolutevalue{\omega_{L,i}\pm\omega_{L,j}}$ | ($+$ for squeezing, -- for beamsplitter couplings). Then, the steady-state amplitudes of the modes (including also pump modes that do not directly appear in the RWA description) are calculated from a set of nonlinear equations derived from the Langevin equations. By linearizing the Langevin equations around this steady state, one can express the quadratic coupling $g_{ij}$ and $\nu_{ij}$ in terms of the bare nonlinear couplings and the steady amplitudes and calculate target values for the latter quantities. In turn, the steady amplitudes can be adjusted by varying the powers and phases of the pumps. The appropriate values of the powers and phases can be estimated from the nonlinear equations for the steady state and can also be fine-tuned directly in the experiments. For most scattering devices, it is possible to achieve the ideal behaviour without taking any asymptotic limit. In this scenario, the cooperativities are typically of order $1$. Asymptotic solutions require large cooperativities. Such cooperativities can be reached in a variety of different hardware platforms featuring optical, microwave, and mechanical modes.
IX Conclusion
Our automated discovery algorithm AutoScatter finds ideal high-level designs for real-world scattering devices like transducers, amplifiers, isolators, or circulators. It provides interpretable and transferable solutions thanks to its graph representation, which encodes in the same set of interpretable weights a whole class of setups, unified by their scattering behavior but with different physical parameters. The flexibility in translating each solution back into physical parameters, allows the user to implement them in a range of state-of-the-art experimental platforms and transfer them to different dissipative regimes, covering both Hamiltonian and reservoir engineering schemes. AutoScatter can discover types of solutions that are normally difficult to obtain in an automated approach, such as asymptotic solutions. Through a series of examples, we have shown that the solutions generated by AutoScatter lead to new insights, in the form of counterintuitive design principles, generalizable design patterns, and analytical solutions to continuous classes of problems.
@@ IX Conclusion
AutoScatter is amenable to numerous future extensions, such as searching for optimized implementations in specific hardware platforms, neural-network-guided heuristics for larger devices where complete enumeration is infeasible, or targeting specific quantum noise characteristics. More generally, our results now open the door towards automated discovery in additional related domains of importance, like photonic transport through periodic structures composed of small modular blocks or scattering-based sensing experiments, as well as transfer to areas like electronic scattering transport. Nonlinear setups or even devices with functionalities based on time-dependent control require significant new developments but still could be treated with the same graph-based efficient search algorithm. Overall, the results shown here validate the fruitfulness of the main ideas guiding the field of artificial scientific discovery.
Reproducibility
We open-source all code used to produce the results of this manuscript in the GitHub repository AutoScatter [9].
@@ Acknowledgements.
We acknowledge fruitful discussions with Mario Krenn, Clara Wanjura, and Flore Kunst. All authors contributed to the ideas, their implementation, and the writing of the manuscript. The implementation of the discovery algorithm and the numerics were performed by J. L.
Appendix A Details of the discrete optimisation
During the optimisation we keep two libraries, one for valid, and one for invalid graphs. At the beginning of the discrete optimization, our method tests only fully connected graphs for validity using the continuous optimisation starting from the graph with zero auxiliary modes. It increments the number of auxiliary modes until it finds the minimum number of auxiliary modes required to achieve the target behavior. This is the smallest fully connected graph that satisfies the target behavior. From there, we investigate which couplings or coupling phases can be removed. This search space corresponds to a directed acyclic graph (see visualization in Fig. 2(b)), where the previously identified fully connected graph is the root node. To search through this space, we perform a breadth-first search. This means that our search always explores all nodes in the search space at a the current depth before proceeding on to the next level starting from the root node.
In each iteration, we first check if the selected graph (the next node in the search space) is already in one of the two libraries. If so, the graph is skipped. Otherwise, it is tested by the continuous optimization. If no solution is found, the tested graph and all its subgraphs are added to the library of invalid graphs. Otherwise, the graph and all its extensions are added to the library of valid graphs. If some of the discovered parameters turn out to be zero, the corresponding simpler graph and all its extensions are added instead. The optimization is finished when all possible graphs have been sorted into one of the two libraries.
Appendix B Details of the continuous optimisation
For the minimization of the loss function in Eq. 3, we use the Broyden–Fletcher–Goldfarb–Shanno algorithm [55], an iterative gradient-descent method for solving non-linear optimization problems. A valid solution is obtained whenever the loss reaches zero. The algorithm fails to find a solution if the loss gets stuck in a non-zero minimum. Since the optimization landscape supports local minima, there is a risk of false-negative errors. To reduce this risk, we repeat the continuous optimization $N_{\rm rep}$ times. The hyperparameter $N_{\rm rep}$ is chosen empirically with a default value $N_{\rm rep}=10$.
During the discrete optimisation we assume that there always exist a valid graph. However, in our work we identified two occasions, where the loss function cannot be minimized to zero for any graph: (i) The desired target scattering behaviour violates fundamental physics. For example, if you ask the noise to adopt a value below the quantum limit, no valid graph can be found. (ii) The desired scattering behaviour can only be realised in a limiting case not covered by our optimisation. If you ask our optimisation scheme to search for the circulator behaviour discussed in Fig. 6 without allowing for far-detuned modes, our continuous optimization will label any graph as invalid.
Please note, that we have assumed in the main text that each port mode is coupled in reflection to a single waveguide. In the Supplemental Material we discuss how some alternative coupling schemes can be translated to our setups and how one can extend our approach to further coupling schemes, covering also setups including running-waves modes that are not invariant under time-reversal, as in [44].
Appendix C Bogoliubov-de Gennes Hamiltonian
@@
Appendix D Rotating Frame
The Hamiltonian in Eq. 5 is defined in a frame in which the phase space of each mode $j$ rotates at the corresponding carrier frequency $\omega_{L,j}$. Thus, $-g_{jj}$ is to be interpreted as the detuning $\Delta_{j}$ of the carrier frequency $\omega_{L,j}$ from the mode resonance $\omega_{j}$, $\Delta_{j}=\omega_{L,j}-\omega_{j}$. For multi-mode circuits, such time-independent Hamiltonian is typically referred to as the rotating-wave Hamiltonian as it is obtained after a Rotating-Wave Approximation (RWA) that drops the fast-oscillating terms, which result from non-resonant interactions. The corrections to the RWA approximation are small if all resonant frequencies $\omega_{j}$ are much larger than the coupling rates $g_{ij}$ and $\nu_{ij}$ as well as the decay rates $\kappa_{j}$.
We note that if all couplings are active the carrier frequencies $\omega_{L,j}$ for the different modes can be chosen independently of each other, adjusting the frequencies $\omega_{ij}^{P}$ of the pump mediating the interaction such that $\omega_{ij}^{P}=\absolutevalue{\omega_{L,i}\pm\omega_{L,j}}$ | ($+$ for squeezing, -- for beamsplitter couplings). Therefore, in experimental scenarios in which the mode resonances $\omega_{j}$ are not easily tunable, one can choose the carrier frequencies to implement the detunings $\Delta_{j}$ required in the scheme. On the other hand, two modes whose coupling is implemented passively (without a parametric pump) have the same carrier frequency and, thus, one of the resonance frequencies together with the common carrier frequency have to be tuned to implement the target detunings.
Besides these constraints, the mode resonances $\omega_{j}$ are irrelevant in the high-level description based on the RWA. This is an appealing feature that makes any insights gained using our method highly transferable across a variety of platforms. A more refined description including non-resonant interactions is typically used to quantify how a specific implementation deviates from the ideal target behavior. This deviation is device-specific and, thus, goes beyond the scope of our work.
Appendix E Langevin equations
@@
$\begin{split}\dot{\hat{a}}_{j}(t)=&-\mathrm{i}\sum_{k}g_{jk}\hat{a}_{k}(t)-\mathrm{i}\sum_{k}\nu_{jk}\hat{a}_{k}^{\dagger}(t)\\ &-\frac{\kappa_{j}+\Gamma_{j}}{2}\hat{a}_{j}(t)-\sqrt{\kappa}_{j}\hat{a}_{j}^{\mathrm{in}}(t)-\sqrt{\Gamma}_{j}\hat{a}_{j}^{\mathrm{noise}}(t)\end{split}$ (7)
Here, $\hat{a}_{j}^{\mathrm{in}}$ is the input field for the main decay channel (for the port modes the out-coupling channel), with decay rate $\kappa_{j}$. Likewise, $\hat{a}_{j}^{\mathrm{noise}}$ is the noise entering port mode $j$ from its intrinsic loss channel, which has decay rate $\Gamma_{j}$. The Langevin equations can be rewritten in a compact form as
$\begin{split}\dot{\hat{\boldsymbol{\xi}}}(t)=&\left(-i\sigma_{z}H_{\rm BdG}-\frac{\kappa+\Gamma}{2}\right)\hat{\boldsymbol{\xi}}(t)\\ &-\sqrt{\kappa}\hat{\boldsymbol{\xi}}^{\mathrm{in}}(t)-\sqrt{\Gamma}\hat{\boldsymbol{\xi}}^{\mathrm{noise}}(t)\end{split}$ (8)